OpenStudy (popgoesthweasal123):

Solve for w. −(1/4w+8)+8=3

OpenStudy (ninja4535):

What do you think is the answer?

OpenStudy (ninja4535):

I can't give you the answer but I can help you solve the problem.

OpenStudy (ninja4535):

I'll just write an example problem and solve it, since you are not answering. So..just follow how I solved the equation. -(2/5y - 17)-17=20 Since there is a negative sign (-) before the parentheses you multiply the number inside the parentheses by 1 That gives you -2/5y +17 -17 = 20 -2/5y +17 -17 =20 -2/5y = 20 Divide 20 by -2/5 Y= -50

OpenStudy (ninja4535):

Wait. Small typo. Multiply the numbers inside the parentheses by -1.

OpenStudy (ninja4535):

I hope this helps.